OpenStudy (anonymous):

Line CD passes through points C(3, –5) and D(6, 0). What is the equation of line CD in standard form?

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ 3}}\quad ,&{\color{blue}{ -5}})\quad &({\color{red}{ 6}}\quad ,&{\color{blue}{ 0}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}} \\ \quad \\ y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values and move the "x" and "y" to the left}\\ \qquad \uparrow\\ \textit{point-slope form}\)

OpenStudy (anonymous):

We use the following fact: \[ \frac{\Delta y}{\Delta x} \]is equal for all points.

OpenStudy (anonymous):

\[ \frac{y-(-5)}{x-3} = \frac{0-(-5)}{6-3} \]

OpenStudy (anonymous):

@jdoe0001 How did you type that all out so quick?

OpenStudy (anonymous):

a) 5x + 3y = 18 b) 5x – 3y = 30 c) 5x – y = 30 d) 5x + y = 18